| Friend | Site | Result | Problem | Text |
|---|---|---|---|---|
| Monari | Medicina | Success | None | None |
| K. Borkowski | Torun | Success | None | No known problems. |
| Michael Lindqvist | Onsala | Minor failures | None | BBC3 did not have the correct filter. Changed around 06.50 UTC. |
| Jonathan Quick | Hartebeesthoek | Minor failures | RFI | Intermittent RFI throughout experiment. |
| Michael Lindqvist | Onsala | Minor failures | RECE | We had to use warm HEMTs |
| Alastair Gunn | Jodrell_Bank | Success | None | The Lovell Telescope was used. No known problems. |
| Alastair Gunn | Cambridge | Failure | LATE | Controller forgot to start Cambridge schedule on time. Did not start until scan 38. No other known problems. |
| Tony Foley | Westerbork | Minor failures | WIND HIGH RFI | Wb used in phased array mode. Still LOTS of snow on the dishes. Calibration poor and Tsys very high. Some very bad RFI |
| A. Kraus | Effelsberg | Success | None | None |